OpenStudy (anonymous):

A jacket originally sold for $45.00. This week, it is on sale for 20% off. What is the cost of the jacket on sale? SHOW WORK

OpenStudy (anonymous):

45*.8 =36 or 45 *.20=9 45-9=36

OpenStudy (anonymous):

sale price = 45.00 - (45.00/100)*20 sale price = 45.00-9 sale price = 36

OpenStudy (anonymous):

because off so still $45

OpenStudy (anonymous):

cost was $45, on sale $36

OpenStudy (anonymous):

\[45-(45*0.2)=\]\[45-9=36\]

jhonyy9 (jhonyy9):

- for your understanding 20% from 45 you can calculate from this if you write it hence : from 100 ..... 20 from 45 ...... x ------------- (45)(20) 900 x= ------- = ---- = 9 100 100 - hence you get the price with the sale for 20% off equal 45-9=36 - hope that you have understood it hence clearly